# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v2

@@ 3.1.3 Feed-Forward-based Methods
Inspired by 2D editing methods [128, 129], editing works based on 3D Gaussian Splatting (3DGS) have advanced. These methods leverage the rich scene representation of 3DGS and often incorporate 2D diffusion models to guide precise 3D modifications. This section reviews and categorizes editing techniques and tasks.
@@ 3.2.1 Editing Based on Text Prompt
$\bullet$ Localizing the Editing Object. Most current research first requires text to locate the objects that need to be edited. GaussianEditor [8, 35], as pioneering work in the field of 3D editing based on 3D Gaussian Splatting technology, proposes innovative editing methodologies. The former significantly enhances editing precision by integrating textual descriptions throughout the training process to track editing targets. The latter achieves refined control over the editing process by extracting regions of interest (ROIs) corresponding to textual instructions and mapping these regions to 3D Gaussians. These methods leverage the InstructPix2Pix [128] model during the editing process, but due to its inherent limitations, issues like inaccurate localization and limited editing control still persist. To address these, GSEditPro [133] introduces an attention-based progressive localization module, leveraging cross-attention layers from the T2I model to classify Gaussians and precisely localize editing regions. It also incorporates DreamBooth [134] for enhanced diversity and flexibility. Xiao et al. [135] propose a lighting-aware 3D scene editing pipeline, utilizing an Anchor View Proposal (AVP) algorithm to identify target regions and a coarse-to-fine optimization process with Depth-guided Inpainting Score Distillation Sampling (DI-SDS) for texture and lighting consistency.
@@
$\bullet$ Efforts toward Multi-View Consistency. Editing 3D objects while maintaining multi-view consistency is a core challenge in the field of 3D editing. GaussCtrl [136] leverages depth-conditioned editing by utilizing naturally consistent depth maps to ensure geometric consistency across multi-view images, along with attention-based latent code alignment to achieve a unified appearance across views. Luo et al. [137] propose a progressive 3D editing strategy using a Trajectory-Anchored Scheme (TAS) and a dual-branch editing mechanism to ensure multi-view consistency, addressing challenges in error accumulation during text-to-image processes. Gomel et al. [138] introduce a geometry-guided warping mechanism that leverages the scene’s depth and structural information to accurately map edits across views, ensuring multi-view consistency. DGE [139] employs spatiotemporal attention to jointly edit the selected key views and inject their features into other views, using correspondences derived from epipolar-constrained visual features, thereby establishing multi-view consistency. SplatFlow [140] leverages a multi-view rectified flow (RF) model to simultaneously generate multi-view images, depth information, and camera poses based on text prompts. By incorporating training-free inversion and inpainting techniques, it achieves seamless editing of 3DGS. InterGSEdit [141] introduces a 3D Geometry-Consistent Attention Prior and an Adaptive Cross-Dimensional Attention Fusion Network to ensure multi-view consistency and enable fine-grained detail recovery.
@@ 3.2.3 Style Transfer
$\bullet$ Efforts toward Objective Function. SGSST [156] designs a Simultaneously Optimized Scales loss, enabling style transfer across all scales for consistent multi-scale stylization. ReGS [157] improves stylization consistency via a Stylized Pseudo View Supervision loss for uniform appearance across views, and a Template Correspondence Matching loss to propagate style to occluded regions. WaSt-3D [158] performs style transfer directly on 3D Gaussians, using an entropy-regularized Wasserstein-2 distance to smoothly map the style scene distribution to the content scene via gradient flow, and decomposing stylization into smaller subproblems for efficiency. Multi-StyleGS [159] introduces a style loss with bipartite matching between multiple style image regions and GS points to enable local style transfer.
@@ 3.2.4 Other Editing Tasks
$\bullet$ Inpainting. InFusion [173] uses an image-conditioned depth completion model to guide point initialization and directly recovers depth maps from images. RefFusion [174] fine-tunes an image inpainting diffusion model, effectively aligning the prior distribution with the 3D target scene. This reduces score distillation variance and yields clearer details.
@@
While NeRF-based methods produce high-quality 3D content, they are slow and computationally expensive. Recent advances in 3D Gaussian Splatting (3DGS) enable faster, more efficient 3D generation from text or images, categorized by the type of output.
@@ 3.3.1 Object-level Generation
A. Per-Scene Optimization Methods: Optimization-based methods extract diffusion priors from powerful 2D foundation diffusion models to guide and optimize 3D representations. DreamFusion [78] first introduces the Score Distillation Sampling (SDS) loss, which optimizes NeRF [2] using images generated from text prompts. Building on this, many methods [175, 176, 177, 178] have extensively explored optimization-based approaches.
@@
$\bullet$ Efforts toward Improving SDS. Although SDS provides guidance for 3D generation, it often suffers from low fidelity and over-smoothing. Recent works address these issues by enhancing geometric accuracy, detail preservation, and stability. (a) Text-to-3D. StableDreamer [187] formulates the SDS prior and L2 reconstruction loss as equivalent, reducing multi-face artifacts and improving geometry. Li et al. [188] propose Guided Consistency Sampling integrated with 3DGS to enhance detail and fidelity. DreamMapping [189] introduces Variational Distribution Mapping, accelerating distribution modeling by treating rendered images as degraded diffusion outputs. HumanGaussian [190] designs a structure-aware SDS for joint optimization of human appearance and geometry. LucidDreamer [41] mitigates over-smoothing via Interval Score Matching (ISM) with deterministic diffusion trajectories, while TSM [191] extends ISM by generating two reverse DDIM [192] paths to reduce cumulative errors and pseudo-ground truth inconsistencies. GaussianMotion [193] employs adaptive score distillation to balance realism and smoothness. (b) Image-to-3D. Basak et al. [194] introduce a frequency-based distillation loss, extracting low-frequency geometry from 3D diffusion priors and refining high-frequency texture with 2D diffusion. GECO [195] adopts a two-stage pipeline combining Variational Score Distillation [196] and multi-view consistency refinement, enabling sub-second high-quality generation. DreamPhysics [197] leverages video diffusion priors to learn material properties and applies motion distillation sampling to emphasize dynamic cues.
@@
$\bullet$ Network Design. Designing feed-forward networks tailored to 3DGS is an active research direction. (a) Text-to-3D. BrightDreamer[214] proposes a fast feed-forward framework combining a text-guided shape deformation network with a triplane generator to efficiently predict and optimize Gaussian attributes. (b) Image-to-3D. Lu et al. [215] present the Point-to-Gaussian Generator composed of multiple APP blocks, each with a point feature extractor, projection module, and cross-modal attention, followed by multi-linear heads to decode 3D Gaussians. GRM [38] introduces an upsampler using a windowed self-attention variant to capture non-local cues and enhance high-frequency details. UniGS [216] develops a DETR-like framework treating Gaussians as queries, updated via multi-view cross-attention over input images to mitigate ghosting artifacts.
$\bullet$ Optimization Diffusion Models. These methods directly produce 3D Gaussians by fine-tuning or training diffusion models. (a) Text-to-3D. GVGEN[204] adopts a structured Gaussian volume representation and a coarse-to-fine pipeline, using a candidate pool strategy for pruning and densification to reconstruct high-fidelity 3D scenes from complex text. (b) Image-to-3D. NovelGS [217] generates 3D Gaussians from sparse-view images with novel-view denoising, achieving state-of-the-art reconstruction with consistent, sharp textures. Ouroboros3D [218] jointly trains multi-view image generation and 3D reconstruction in a recursive diffusion process, enabling mutual adaptation for robust inference. Cycle3D [219] cyclically integrates 2D diffusion-based generation and feed-forward 3D reconstruction to improve multi-view consistency and texture quality. DiffusionGS [220] outputs Gaussian point clouds at each timestep for view-consistent generation from any direction, introducing a scene–object mixed training strategy to enhance scalability. Hi3D [221] employs a video diffusion paradigm with 3D-aware priors and video-to-video refinement for consistent multi-view reconstruction.
@@ 3.3.2 Scene-level Generation
B. Iterative Generation Methods: These approaches iteratively refine 3D Gaussian representations through repeated rendering and optimization to achieve high-quality scene reconstruction. (a) Text-to-3D. Text2Room [42] combines monocular depth estimation with a text-conditioned inpainting model, aligning scene frames with existing geometry for 3D reconstruction. Text2Immersion [227] progressively generates Gaussian point clouds using 2D diffusion and depth models, followed by refinement via interpolation. RealmDreamer [228] initializes points with a text-to-image generator, lifts them into 3D, computes occlusion volumes, and optimizes across multiple views using an image-conditioned diffusion model. WonderJourney [229] employs LLMs to generate scene descriptions, builds 3D scenes via text-driven point cloud generation, and validates them with VLMs. HoloDreamer [230] addresses global inconsistency and incompleteness by generating high-resolution panoramas for holistic initialization, followed by rapid 3DGS-based reconstruction for view-consistent, enclosed scenes. (b) Image-to-3D. LucidDreamer [231] alternates between Dreaming, which generates multi-view consistent images guided by a point cloud, and Alignment, which integrates new 3D points into a unified scene, producing realistic 3D Gaussian splats from text, RGB, or RGBD inputs. VistaDream [232] enforces multi-view consistency during reverse diffusion sampling to improve temporal coherence. Kang et al. [233] propose a transformer-based latent diffusion model with explicit view geometry constraints, incorporating warped feature maps, epipolar-weighted source features, Plücker ray maps, and camera poses for high-quality novel view generation.
@@
$\bullet$ Network Design. (a) Text-to-3D. TextSplat [234] introduces Text-Guided Semantic Fusion Module to integrate multi-source semantic features under sentence-level guidance, enhancing geometry semantic consistency. (b) Image-to-3D. PixelSplat [235] employs a multi-view epipolar transformer to address scale ambiguity, while MVSplat [4] and Splatter360 [236] construct planar or spherical cost volumes for improved geometry estimation. MVSplat360 [237] leverages pre-trained video diffusion for 3D-consistent views. GaussianCity [238] adopts BEV-Point as a compact intermediate representation with a spatial-aware decoder. SelfSplat [239] introduces matching-aware pose estimation and depth refinement to ensure geometric consistency. CATSplat [240] fuses vision–language text features with point cloud features via transformer blocks, and OmniSplat [241] uses a Yin–Yang grid to reduce distortion.
$\bullet$ Optimization Diffusion Models. (a) Text-to-3D. Prometheus[242] sequentially trains GS-VAE and MV-LDM for direct text-to-3D scene generation. VideoRFSplat [243] couples a pose generation model with a pretrained video generator via communication blocks to produce multi-view images and camera poses. (b) Image-to-3D. ViewCrafter [244] employs a point-conditioned video diffusion model for high-quality 3D scenes, while Wonderland [245] integrates diverse camera trajectories via a dual-branch conditioning mechanism for view-consistent latents. GGS [246] combines 3D representations with latent video diffusion and a custom decoder to synthesize scenes from feature fields. VideoScene [247] distills video diffusion through a 3D-aware leap flow strategy for one-step generation. SceneSplatter [248] adopts a momentum-based paradigm to balance generative priors and existing scene information during generation.
3.4 Other Application Tasks
3.4.1 Human Avatar
3.4.2 SLAM
3.4.3 3DGS Object Detection
3D object detection focuses on identifying and localizing objects in 3D space. Similar to segmentation, detection methods can be categorized into prompt-based and generic approaches. Prompt-based detection, often referred to as open-vocabulary detection or visual grounding, takes textual inputs (e.g., category names or language expressions) and outputs object locations. Most 3DGS-based segmentation methods also support this type of detection by generating the maximum response point, and are thus not detailed here. Specifically, SpatialReasoner [275] combines LLM-driven spatial reasoning with a hierarchical feature field enhanced by visual properties. It distills CLIP and SAM features and leverages fine-tuned LLMs to infer spatial instructions, enabling precise localization of target instances based on relational language cues. For generic object detection, representative methods include: Gaussian-Det [276], which models objects continuously using input Gaussians as feature descriptors across partial surfaces; 3DGS-DET [277], which incorporates boundary guidance and box-focused sampling to refine Gaussian distributions and suppress background noise; and MATT-GS [278], which employs masked attention 3D Gaussian Splatting to enhance object localization.
[63][26][279][7][22][280][75][281][67][67][18][282][283][284]
@@ 4 Performance Comparison
TABLE I: Statistics of representative 3DGS segmentation datasets. See §4.1 for more detailed descriptions.
  Dataset Venue #Scene #Views(Avg.) Characterization
TABLE II: Quantitative 3D instance segmentation on Replica[285] and LERF-Mask[22]. Following [20], we report mIoU and F-score on Replica, and mIoU and mBIoU on LERF-Mask.

### v3

@@ 3.1.4 Related Task: 3DGS Object Detection
Inspired by 2D editing methods [141, 142], editing works based on 3D Gaussian Splatting (3DGS) have advanced. These methods leverage the rich scene representation of 3DGS and often incorporate 2D diffusion models to guide precise 3D modifications. This section reviews and categorizes editing techniques and tasks, with an illustrative overview provided in Fig. 4.
@@ 3.2.1 Editing Based on Text Prompt
$\bullet$ Localizing the Editing Object. Most current research first requires text to locate the objects that need to be edited. GaussianEditor [8, 35], as pioneering work in the field of 3D editing based on 3D Gaussian Splatting technology, proposes innovative editing methodologies. The former significantly enhances editing precision by integrating textual descriptions throughout the training process to track editing targets. The latter achieves refined control over the editing process by extracting regions of interest (ROIs) corresponding to textual instructions and mapping these regions to 3D Gaussians. These methods leverage the InstructPix2Pix [141] model during the editing process, but due to its inherent limitations, issues like inaccurate localization and limited editing control still persist. To address these, GSEditPro [146] introduces an attention-based progressive localization module, leveraging cross-attention layers from the T2I model to classify Gaussians and precisely localize editing regions. It also incorporates DreamBooth [147] for enhanced diversity and flexibility. Xiao et al. [148] propose a lighting-aware 3D scene editing pipeline, utilizing an Anchor View Proposal (AVP) algorithm to identify target regions and a coarse-to-fine optimization process with Depth-guided Inpainting Score Distillation Sampling (DI-SDS) for texture and lighting consistency. RoMaP [149] employs 3D-Geometry Aware Label Prediction to identify editing targets and combines SDS loss with additional regularization to produce high-quality local edits while preserving contextual consistency.
@@
$\bullet$ Efforts toward Multi-View Consistency. Editing 3D objects while maintaining multi-view consistency is a core challenge in the field of 3D editing. GaussCtrl [150] leverages depth-conditioned editing by utilizing naturally consistent depth maps to ensure geometric consistency across multi-view images, along with attention-based latent code alignment to achieve a unified appearance across views. Luo et al. [151] propose a progressive 3D editing strategy using a Trajectory-Anchored Scheme (TAS) and a dual-branch editing mechanism to ensure multi-view consistency, addressing challenges in error accumulation during text-to-image processes. Gomel et al. [152] introduce a geometry-guided warping mechanism that leverages the scene’s depth and structural information to accurately map edits across views, ensuring multi-view consistency. DGE [153] employs spatiotemporal attention to jointly edit the selected key views and inject their features into other views, using correspondences derived from epipolar-constrained visual features, thereby establishing multi-view consistency. SplatFlow [154] leverages a multi-view rectified flow (RF) model to simultaneously generate multi-view images, depth information, and camera poses based on text prompts. By incorporating training-free inversion and inpainting techniques, it achieves seamless editing of 3DGS. InterGSEdit [155] introduces a 3D Geometry-Consistent Attention Prior and an Adaptive Cross-Dimensional Attention Fusion Network to ensure multi-view consistency and enable fine-grained detail recovery. Pro3D-Editor [156] introduces Primary-view Sampler and Key-view Render to prevent inconsistency conflicts during multi-view editing. DFFSplat [157] integrates a 3D-consistent diffusion feature field into the editing pipeline and uses a dual-encoder architecture to disentangle view-independent structure from view-dependent appearance, enabling consistent multi-view editing.
@@ 3.2.3 Style Transfer
$\bullet$ Efforts toward Objective Function. SGSST [114] designs a Simultaneously Optimized Scales loss, enabling style transfer across all scales for consistent multi-scale stylization. ReGS [171] improves stylization consistency via a Stylized Pseudo View Supervision loss for uniform appearance across views, and a Template Correspondence Matching loss to propagate style to occluded regions. WaSt-3D [172] performs style transfer directly on 3D Gaussians, using an entropy-regularized Wasserstein-2 distance to smoothly map the style scene distribution to the content scene via gradient flow, and decomposing stylization into smaller subproblems for efficiency. Multi-StyleGS [173] introduces a style loss with bipartite matching between multiple style image regions and GS points to enable local style transfer. CLIPGaussian [174] jointly leverages CLIP and VGG features to enable style transfer while preserving the original content.
@@
$\bullet$ Feed-Forward-based Model. Styl3R [183] introduces a branched architecture that separates structural modeling from appearance shading, effectively preventing style transfer from distorting the underlying 3D scene structure, while employing an identity loss to facilitate training of the stylized model. $\text{A}^{3}$GS [184], a novel feed-forward framework, introduces an AdaIN-based stylizer that effectively injects the style features of a target image into the latent representations of a 3DGS scene.
@@ 3.2.4 Other Editing Tasks
$\bullet$ Inpainting. InFusion [188] uses an image-conditioned depth completion model to guide point initialization and directly recovers depth maps from images. RefFusion [189] fine-tunes an image inpainting diffusion model, effectively aligning the prior distribution with the 3D target scene. This reduces score distillation variance and yields clearer details. PAInpainter [190] integrates inpainting view sampling, cross-view content propagation, and consistency verification to enable 3D Gaussian inpainting.
@@
While NeRF-based methods produce high-quality 3D content, they are slow and computationally expensive. Recent advances in 3D Gaussian Splatting (3DGS) enable faster, more efficient 3D generation from text or images, categorized by the type of output. Figure 5 provides representative generation pipelines.
@@ 3.3.1 Object-level Generation
A. Per-Instance Optimization Methods: Optimization-based methods extract diffusion priors from powerful 2D foundation diffusion models to guide and optimize 3D representations. DreamFusion [78] first introduces the Score Distillation Sampling (SDS) loss, which optimizes NeRF [2] using images generated from text prompts. Building on this, many methods [191, 192, 193, 194] have extensively explored optimization-based approaches.
@@
Figure 5: A comparison of four 3DGS generation pipelines. The examples are adapted from LucidDreamer [41], Text2Room [42], LGM [37], and CATSplat [203].
$\bullet$ Efforts toward Improving SDS. Although SDS provides guidance for 3D generation, it often suffers from low fidelity and over-smoothing. Recent works address these issues by enhancing geometric accuracy, detail preservation, and stability. (a) Text-to-3D. StableDreamer [204] formulates the SDS prior and L2 reconstruction loss as equivalent, reducing multi-face artifacts and improving geometry. Li et al. [205] propose Guided Consistency Sampling integrated with 3DGS to enhance detail and fidelity. DreamMapping [206] introduces Variational Distribution Mapping, accelerating distribution modeling by treating rendered images as degraded diffusion outputs. HumanGaussian [207] designs a structure-aware SDS for joint optimization of human appearance and geometry. LucidDreamer [41] mitigates over-smoothing via Interval Score Matching (ISM) with deterministic diffusion trajectories, while TSM [208] extends ISM by generating two reverse DDIM [209] paths to reduce cumulative errors and pseudo-ground truth inconsistencies. GaussianMotion [210] employs adaptive score distillation to balance realism and smoothness. (b) Image-to-3D. Basak et al. [211] introduce a frequency-based distillation loss, extracting low-frequency geometry from 3D diffusion priors and refining high-frequency texture with 2D diffusion. GECO [212] adopts a two-stage pipeline combining Variational Score Distillation [213] and multi-view consistency refinement, enabling sub-second high-quality generation. DreamPhysics [214] leverages video diffusion priors to learn material properties and applies motion distillation sampling to emphasize dynamic cues. TraCe [215] reformulates the Schrödinger Bridge into an explicit diffusion bridge linking the current rendering to its text-conditioned, denoised target, while training a LoRA-adapted model on the trajectory’s score dynamics for robust 3D optimization.
@@
$\bullet$ Network Design. Designing feed-forward networks tailored to 3DGS is an active research direction. (a) Text-to-3D. BrightDreamer [232] proposes a feed-forward framework combining a text-guided shape deformation network with a triplane generator to predict and optimize Gaussian attributes. (b) Image-to-3D. Lu et al. [233] present the Point-to-Gaussian Generator composed of multiple APP blocks, each with a point feature extractor, projection module, and cross-modal attention, followed by multi-linear heads to decode 3D Gaussians. GRM [38] introduces an upsampler using a windowed self-attention to capture non-local cues and enhance high-frequency details. UniGS [234] is a DETR-like model treating Gaussians as queries, updated via multi-view cross-attention over input images to reduce ghosting artifacts.
$\bullet$ Optimization Diffusion Models. These methods directly produce 3D Gaussians by fine-tuning or training diffusion models. (a) Text-to-3D. GVGEN[222] adopts a structured Gaussian volume representation and a coarse-to-fine pipeline, using a candidate pool strategy for pruning and densification to reconstruct high-fidelity 3D scenes from complex text. (b) Image-to-3D. NovelGS [235] generates 3D Gaussians from sparse-view images with novel-view denoising, achieving state-of-the-art reconstruction with consistent, sharp textures. Ouroboros3D [236] jointly trains multi-view image generation and 3D reconstruction in a recursive diffusion process, enabling mutual adaptation for robust inference. Cycle3D [237] cyclically integrates 2D diffusion-based generation and feed-forward 3D reconstruction to improve multi-view consistency and texture quality. DiffusionGS [238] outputs Gaussian point clouds at each timestep for view-consistent generation from any direction, introducing a scene–object mixed training strategy to enhance scalability. Hi3D [239] employs a video diffusion paradigm with 3D-aware priors and video-to-video refinement for consistent multi-view reconstruction. TRIM [240] employs a lightweight selector model to assess latent Gaussian primitives from multiple noise samples and prunes early trajectories by selecting high-quality candidates, thereby accelerating inference. GSV3D [241] utilizes gaussian splatting for geometric distillation within the SVD [242] framework, ensuring structural consistency to boost 3D generation quality.
[63][26][243][7][22][244][75][245][67][67][18][246][247][248]
@@ 3.3.2 Scene-level Generation
B. Iterative Generation Methods: These approaches iteratively refine 3D Gaussian representations through repeated rendering and optimization to achieve high-quality scene reconstruction. (a) Text-to-3D. Text2Room [42] combines monocular depth estimation with a text-conditioned inpainting model, aligning scene frames with existing geometry for 3D reconstruction. Text2Immersion [254] progressively generates Gaussian point clouds using 2D diffusion and depth models, followed by refinement via interpolation. RealmDreamer [255] initializes points with a text-to-image generator, lifts them into 3D, computes occlusion volumes, and optimizes across multiple views using an image-conditioned diffusion model. WonderJourney [256] employs LLMs to generate scene descriptions, builds 3D scenes via text-driven point cloud generation, and validates them with VLMs. HoloDreamer [257] addresses global inconsistency and incompleteness by generating high-resolution panoramas for holistic initialization, followed by rapid 3DGS-based reconstruction for view-consistent, enclosed scenes. (b) Image-to-3D. ViewCrafter [258] employs a point-conditioned video diffusion model for high-quality 3D scenes. LucidDreamer [259] alternates between Dreaming, which generates multi-view consistent images guided by a point cloud, and Alignment, which integrates new 3D points into a unified scene, producing realistic 3D Gaussian splats from text, RGB, or RGBD inputs. VistaDream [260] enforces multi-view consistency during reverse diffusion sampling to improve temporal coherence. Kang et al. [261] propose a transformer-based latent diffusion model with explicit view geometry constraints, incorporating warped feature maps, epipolar-weighted source features, Plücker ray maps, and camera poses.
@@
$\bullet$ Network Design. (a) Text-to-3D. TextSplat [262] introduces Text-Guided Semantic Fusion Module to integrate multi-source semantic features under sentence-level guidance, enhancing geometry semantic consistency. (b) Image-to-3D. PixelSplat [263] employs a multi-view epipolar transformer to address scale ambiguity, while MVSplat [4] and Splatter360 [264] construct planar or spherical cost volumes for improved geometry estimation. MVSplat360 [265] leverages pre-trained video diffusion for 3D-consistent views. GaussianCity [266] adopts BEV-Point as a compact intermediate representation with a spatial-aware decoder. SelfSplat [267] introduces matching-aware pose estimation and depth refinement to ensure geometric consistency. CATSplat [203] fuses vision–language text features with point cloud features, and OmniSplat [268] uses a Yin–Yang grid to reduce distortion.
$\bullet$ Optimization Diffusion Models. (a) Text-to-3D. Prometheus[269] sequentially trains GS-VAE and MV-LDM for direct text-to-3D scene generation. VideoRFSplat [270] couples a pose generation model with a pretrained video generator via communication blocks to produce multi-view images and camera poses. (b) Image-to-3D. Wonderland [271] integrates diverse camera trajectories via a dual-branch conditioning mechanism for view-consistent latents. GGS [272] combines 3D representations with latent video diffusion and a custom decoder to synthesize scenes from feature fields. VideoScene [273] distills video diffusion through a 3D-aware leap flow strategy for one-step generation. SceneSplatter [274] adopts a momentum-based paradigm to balance generative priors and existing scene information during generation.
@@ 4 Performance Comparison
4.1 Performance Benchmarking: 3DGS Segmentation
$\bullet$ Datasets. Appendix Table A.1 summarizes the key features of the datasets for 3DGS segmentation.
TABLE I: Quantitative 3D instance segmentation on Replica[275] and LERF-Mask[22]. Following [20], we report mIoU and F-score on Replica, and mIoU and mBIoU on LERF-Mask.